It takes Nate five hours to paint a wall, and it takes Jill seven hours to paint the same wall. About how many hours will it will take if Nate and Jill work together to paint the wall?
Answer by stanbon(75887)

It takes Nate five hours to paint a wall, and it takes Jill seven hours to paint the same wall. About how many hours will it will take if Nate and Jill work together to paint the wall?
-------------
rate + rate = together rate
1/5 + 1/7 = 1/x
7x + 5x = 5*7
-----
12x = 35
x = 2.92 hrs = 2 hr 55 minutes (time to do the job together)
----------
Cheers,
Stan H.
----------